What length of conductor, placed at right angles to a magnetic field of magnitude 0.052 T, and with a current of 16.04 A, experiences a force of 1.013 N?




1
Expert's answer
2021-06-02T09:35:07-0400
"F=BILsin\theta,""L=\dfrac{F}{BIsin\theta},""L=\dfrac{1.013\ N}{0.052\ T\cdot16.04\ A\cdot sin90^{\circ}}=1.21\ m."